OpenStudy (anonymous):

The height of a small rocket fired straight upward is measured by the function h(t) = −16t2 + 512t, where h(t) is measured in feet and t is time in seconds. When is the rocket at a height of 960 ft?

OpenStudy (anonymous):

-16t^2 + 512t = 960 -16t^2 + 512t -960 = 0 -t^2 + 32t - 60 = 0 t^2 - 32t + 60 = 0 (t-2)(t-30)=0 t = 2 or 20

OpenStudy (anonymous):

*2 or 30